# MemoFurther Readings

### v1

@@ Abstract
Introduction
@@
Social media has evolved into a vital arena for public discourse, serving as a platform where individuals and communities can converge to discuss political, social, and cultural issues. These platforms have been instrumental in fostering large-scale movements, such as the Arab Spring and Black Lives Matter, where activists and citizens alike have used social media to amplify calls for justice and change (Howard and Hussain, 2013; Freelon et al., 2016). However, social media has also been linked to polarization and radicalization, as exemplified by the events surrounding the U.S. Capitol attack (Election Integrity Partnership, 2020; Luceri et al., 2021), where online platforms played a role in mobilizing and coordinating participants  (Vishnuprasad et al., 2024). Similar cases include the Christchurch mosque shootings (Fakhrurroja et al., 2019) and far-right rallies across Europe (Klein and Muis, 2019; Pérez-Curiel, 2020), where social media content has been shown to exacerbate divisive sentiments and mobilize fringe communities (Davey and Ebner, 2019). The potential for influence of online discourse has contributed to the establishment of state-sponsored information operations aimed at steering public opinion by introducing false or misleading narratives on popular platforms (Nisbet and Kamenchuk, 2019; Wilson et al., 2018; Molter and DiResta, 2020).
@@
Coordinated inauthentic behavior represents a prominent tactic for distorting public discourse by amplifying specific viewpoints and giving the illusion of widespread support (Giglietto et al., 2020; Khaund et al., 2021; Pacheco et al., 2021; Murero, 2023; Cima et al., 2024). These activities typically involve actions such as synchronized posting and co-activity behavior patterns (Oates and Gray, 2019; Pacheco et al., 2021; Luceri et al., 2024a)— such as similar retweets, hashtag sequences, link sharing— designed to push particular narratives to the forefront of public attention. Studies have also documented how these coordinated behaviors contribute to a range of detrimental social effects, such as the spread of propaganda (Hristakieva et al., 2022a), conspiracy theories (Tardelli et al., 2024), and the promotion of disinformation (Weber and Neumann, 2021). Moreover, coordinated online actions have been linked to heightened toxicity in online conversations (Loru et al., 2024) and the dissemination of extremist ideologies (Kursuncu et al., 2021; Vishnuprasad et al., 2024).
@@
In recent years, a growing body of research has focused on the role of social media in the context of political elections. For example, studies on the Brazilian elections have documented the strategic use of bots and coordinated networks to shape public opinion and influence voter perceptions, highlighting the risks of computational propaganda in electoral processes (Pacheco, 2024). Similarly, research on European elections has illustrated how misinformation campaigns spread across social media can alter public perceptions, as seen in countries such as France, Germany, and Italy (Ferrara, 2017; Lovari, 2020; Baqir et al., 2024; Nogara et al., 2024). In the United States, coordinated online activities around elections have garnered significant attention, with investigations revealing how coordinated bot activities and foreign information operations have sought to manipulate voter beliefs and amplify divisive content across multiple election cycles (Bessi and Ferrara, 2016; Howard et al., 2018; Luceri et al., 2019b; Tardelli et al., 2024). Notably, this large body of literature has predominantly focused on identifying such operations within individual platforms.
@@
In this work, we expand on current research in coordination detection by broadening our scope to include the identification of coordinated inauthentic activity (CoIA) spanning multiple platforms. We build on, adapt, and advance state-of-the-art coordination detection techniques to identify intra- and cross-platform CoIA that promotes external web domains and amplify specific narratives in the context of the 2024 U.S. Presidential Election. We leverage a large-scale dataset covering election-related online conversations spanning several platforms, including Twitter/$\mathbb{X}$, Facebook, and Telegram, collected during the run-up months to the Presidential Election. Combining insights from computational techniques for CoIA detection and language models for content analysis, we aim to answer the following Research Questions (RQs):
@@
Leveraging our multi-platform dataset and advancing CoIA detection techniques, we uncover multiple networks of coordinated inauthentic accounts. We analyzed the textual content and link-sharing behaviors of these accounts, identifying the narratives they aim to amplify and the external domains they direct traffic to, assessing both their credibility and ties to foreign and domestic entities. Our findings reveal coordinated efforts to promote Russian-affiliated media across Telegram and $\mathbb{X}$, with highly partisan, low-credibility content systematically amplified by these networks. Conspiracy theories surrounding public health, the environment, and political topics, such as immigration and geopolitical tensions, were prominently featured. Notably, QAnon-related narratives were especially prevalent on Telegram, with coordinated accounts driving much of the discussion. Furthermore, we assessed the prevalence of AI-generated content produced by coordinated actors and the level of engagement their content attracted. We found that coordinated actors on Telegram relied on AI-generated content significantly more than organic users, while the opposite trend was observed on Facebook. This study sheds new light on cross-platform coordination efforts related to the upcoming U.S. Presidential Election, revealing the complex dynamics of influence operations. These findings underscore the urgent need for regulatory measures that go beyond individual platforms to effectively tackle the growing challenge of cross-platform coordinated inauthentic activity.
Related Work
Influence Campaigns Interfering with U.S. Elections
@@
A growing body of research has focused on the influence of coordinated disinformation campaigns during U.S. elections, particularly through the manipulation of social media platforms (Sharma et al., 2022; Ferrara et al., 2020). Studies have uncovered significant efforts by foreign and domestic actors to manipulate the political landscape during both the 2016 and 2020 U.S. Presidential Elections (Bessi and Ferrara, 2016; Vishnuprasad et al., 2024). For example, the Russian Internet Research Agency (IRA) was linked to a large-scale disinformation campaign during the 2016 election, deploying thousands of social bots and state-sponsored trolls to promote divisive narratives and foster discord (Badawy et al., 2019, 2018). The activity of these inauthentic actors has been extensively studied on Twitter (Zannettou et al., 2020; Luceri et al., 2020; Addawood et al., 2019; Luceri et al., 2019a), with fewer studies examining their presence on other platforms (Zannettou et al., 2019b, a).
@@
Similarly, the 2020 U.S. Presidential Election witnessed the propagation of various false claims and conspiracy theories, including narratives about voter fraud and COVID-19 misinformation (Ferrara et al., 2020). Prominent examples include the “Stop the Steal” movement, which spread across platforms like Twitter and Facebook, inciting allegations that the election results were fraudulent (Election Integrity Partnership, 2020). Coordinated activity surrounding this narrative was widespread, often involving amplification techniques such as automated retweets or sharing similar content across accounts to create an illusion of widespread consensus (Luceri et al., 2021).
Detection of Coordinated Inauthentic Activity
@@
Coordinated inauthentic activity has been identified as a predominant tactic in spreading disinformation and conspiratorial content (Luceri et al., 2024a; Pacheco et al., 2020). Researchers have developed sophisticated methods to detect coordination. Coordination is not limited to malicious campaigns: it can also encompass legitimate organizing efforts, social movements, etc. However, in the context of disinformation campaigns, coordinated behavior is typically characterized by manipulative actions aimed at amplifying false narratives (Pacheco et al., 2021; Nizzoli et al., 2021).
@@
The detection of such deceptive, orchestrated efforts has evolved to address both automated and human-coordinated activities through advanced machine learning and network-based approaches. Machine learning techniques have traditionally focused on identifying automated activities, such as bot-like behavior, and distinguishing them from human actions (Yang et al., 2019; Cresci et al., 2016). Recently, attention has shifted toward characterizing suspicious human-operated accounts, with research emphasizing content-, behavioral-, and sequence-based methods to detect coordinated actions by state-sponsored trolls (Hristakieva et al., 2022b; Alizadeh et al., 2020; Luceri et al., 2024, 2020; Nwala et al., 2023; Ezzeddine et al., 2023).
@@
Network-based detection, however, has gained prominence due to its ability to reveal coordinated behavior by constructing networks that highlight similarities in user actions, such as shared content, hashtags, or synchronized posting times (Pacheco et al., 2020; Pacheco et al., 2021; Nizzoli et al., 2021; Mannocci et al., 2024; Magelinski et al., 2022; Luceri et al., 2024a). These networks are analyzed using properties like node centrality and edge weight, which help identify clusters of users involved in coordinated IOs (Luceri et al., 2024a; Vishnuprasad et al., 2024; Tardelli et al., 2023). This approach has proven effective in uncovering both automated and human-coordinated activities, providing critical insights into the tactics and structure of influence campaigns.
@@
In contrast to previous cross-platform studies (Ng et al., 2022; Gatta et al., 2023; Zannettou et al., 2019a), our approach introduces a more advanced and fine-grained coordination detection framework, specifically crafted to minimize false positives by combining network measures (edge weights and node centralities) to identify a highly coordinated set of inauthentic accounts across multiple platforms. Unlike existing methods, which have explored dynamics between mainstream and fringe platforms, our work leverages state-of-the-art language models and AI-generated content detectors to provide a nuanced characterization of users across a broader selection of mainstream social media platforms.
Data
@@
Collection. The rationale behind our data collection is to capture the online discourse surrounding the 2024 U.S. Presidential Election across multiple social media networks. The data collection for our study spans May and June 2024 and covers three major online platforms: Facebook, $\mathbb{X}$ (formerly Twitter), and Telegram. Each platform was queried to obtain posts or messages containing specific election-related keywords. The full list of keywords used to filter and collect data is provided in Appendix 1.
@@
Facebook data was collected through (the now defunct) Crowdtangle, which offered access to public posts of groups and pages. For $\mathbb{X}$, we gather publicly available information, including original tweets, retweets, replies, and quotes, retrieved via the platform’s web interface. Finally, Telegram data was gathered using the Telegram API, allowing the extraction of public chats’ details, meta data, messages, and message attachments. Details on the $\mathbb{X}$ and Telegram data collection infrastructure can be found in (Balasubramanian et al., 2024a; Blas et al., 2024b).
@@
Platform Accounts Posts/Tweets/Messages URLs Domains
@@
Table 1. Datasets statistics.
Statistics. For each platform, the collected content consists of posts on Facebook, tweets on $\mathbb{X}$, and messages on Telegram. To identify potential coordination among accounts operating within each platform, we analyze the textual similarity of their shared content as well as similarity patterns in co-URL sharing. To establish coordination across platforms, we examine only the co-sharing of URLs, as they represent specific pieces of information that can be easily tracked across different platforms (Luceri et al., 2024b). In fact, URLs can be embedded in user’s posts on each platform under analysis. These URLs serve as a common thread linking sharing activities across the Web. Thereby, we identify and count unique URLs embedded within the shared content of each dataset. Table 1 presents a summary of the datasets, including the number of accounts, [^1] We will use the term accounts to also refer to Facebook pages, Facebook groups, and Telegram channels. posts, and unique URLs for each platform.
Methodology
Exposing CoIA within and across Social Media Platforms
@@
Coordinated accounts can execute their campaigns using various strategies; here, we analyze campaigns that focus on promoting specific URLs (or web domains) and campaigns that produce highly similar textual content to amplify certain topics. The former strategy is commonly used to create the illusion of public consensus around certain viewpoints by artificially amplifying links to external webpages, mock websites, and other social media networks (Minici et al., 2024a; Gabriel et al., 2023). The latter is often employed to manipulate platform feed algorithms by pushing specific keywords or hashtags, attempting to boost trending topics, and making content appear more popular than it actually is (Pacheco et al., 2021; Vishnuprasad et al., 2024). We refer to these orchestrated efforts as web domain promotion and content amplification. Since URL-sharing functions through a consistent mechanism across all observed platforms, we analyze web domain promotion both within and across platforms. Due to substantial variability in textual content-sharing—driven by platform-specific constraints, such as character limits on some platforms, and differences in the nature of messaging (e.g., Telegram chats versus Facebook and $\mathbb{X}$ posts)—our analysis of content amplification is focused on within-platform activity.
Detection of Web Domain Promotion
@@
To identify orchestrated campaigns promoting specific web domains, we examine user similarities based on the URLs they share in their social media posts. To ensure high-quality data and reduce noise, we applied several preprocessing steps. First, we imposed a minimum activity threshold, requiring each user to have shared at least 10 unique URLs. This criterion is consistent with prior work (Luceri et al., 2024a; Pacheco et al., 2021) and ensures that the analysis focuses on users with substantial contributions, filtering out those with minimal engagement that could distort the detection of coordinated accounts. Next, we expanded shortened or obfuscated URLs using a URL expansion library [^2] https://github.com/dfreelon/unspooler.
@@
Co-URL similarity network. We constructed a bipartite user-URL network, where users and URLs are connected based on the URLs that users share. In the adjacency matrix of the bipartite graph, users are represented by the rows, whereas the columns represent the URLs. In accordance with previous work (Luceri et al., 2024a; Pacheco et al., 2021), we applied the Term Frequency-Inverse Document Frequency (TF-IDF) transformation to represent the user-URL matrix. To avoid bias towards overly frequent URLs, we set a maximum document frequency (max DF) of the 90th percentile and a minimum document frequency (min DF) threshold of 5 occurrences per URL. This ensures that both very rare and overly common URLs are excluded from the analysis, leaving us with a meaningful set of URLs that can identify CoIAs.
The similarity between users is built by comparing their pairwise co-shared URL patterns. This is obtained by projecting the bipartite network into a user-to-user similarity network. For network construction, we computed exhaustive pairwise cosine similarity between user vectors of the bipartite graph. This similarity measure captures how similar users are based on the URLs they shared, with higher similarity scores indicating stronger coordination. This co-URL similarity network forms the foundation to identify coordinated users potentially driving campaigns aimed at redirecting traffic to specific web domains.
@@
Extraction of network properties for coordinated accounts detection. To identify online coordination, we build upon and integrate state-of-the-art strategies that filter either low-weight edges or peripheral nodes from the similarity network to detect CoIAs. The first approach (Pacheco et al., 2021) prioritizes the strength of similarity, filtering out low-weight edges likely representing spurious similarities. The second approach (Luceri et al., 2024a) focuses on the breadth of similarities, pruning nodes based on their centrality. Luceri et al. (Luceri et al., 2024a) demonstrated that eigenvector centrality is an effective network property for identifying the most suspicious users in a similarity network, where coordinated accounts tend to share numerous similarities with other highly connected nodes. The assumption is that coordinated activity typically involves multiple accounts. In a similarity network, coordinated actions manifest as a pronounced collective similarity, where a coordinated account is highly connected (i.e., similar) to numerous other nodes that are themselves well-connected. Consequently, eigenvector centrality has proven to be an effective network property for detecting online coordination (Luceri et al., 2024a). Here, we extend this notion of collective similarity by considering the density of the similarity network. Network density is a measure that quantifies how close a network is to being fully connected and, in this scenario, it provides an exact measure of collective similarity.
In particular, we use network density as a variable to modulate the combination of thresholds for both node and edge filtering. By integrating these strategies, we aim to leverage the strengths of both filtering techniques, constructing a model that accounts for both the breadth and strength of similarities. This integrated approach is particularly important, as these techniques have not yet been applied to platforms like Telegram and Facebook, where coordination dynamics may vary significantly.
Density-based unsupervised network dismantling. To combine these strategies in a fully unsupervised manner, we developed a novel method for conducting a grid search of parameters in the two-dimensional space defined by node centrality and edge weight distributions. Specifically, the grid search explores two parameters—the quantile of node centrality and the quantile of edge similarity—and evaluates network density for each combination and every connected component in the similarity graph. We hypothesize that, as we progressively filter the similarity graph by removing edges and nodes, a transitional phase in the density of the connected components will signal potentially coordinated accounts. To ensure robustness, we adopt a conservative approach by evaluating the transitional phase of the smallest density among all connected components in the filtered similarity graph. By controlling the component with the lowest density, we ensure that all other components exhibit higher densities. This conservative choice is driven by the unsupervised nature of the task, prioritizing the minimization of false positives, i.e., legitimate users misclassified as coordinated. Therefore, filtering thresholds were chosen based on the transitional phase of the smallest density of a connected component of the similarity graph, thus, focusing on identifying highly suspicious CoIA. Details of the parameters used for detection are provided in the Appendix.
Detection of Content Amplification

### v2

# Exposing Cross-Platform Coordinated Inauthentic Activity in the Run-Up to the 2024 U.S. Election

@@ Abstract.
Information Operations, Coordinated Inauthentic Behavior
journalyear: 2025copyright: ccconference: Proceedings of the ACM Web Conference 2025; April 28-May 2, 2025; Sydney, NSW, Australiabooktitle: Proceedings of the ACM Web Conference 2025 (WWW ’25), April 28-May 2, 2025, Sydney, NSW, Australiadoi: 10.1145/3696410.3714698isbn: 979-8-4007-1274-6/25/04ccs: Human-centered computing Social networksccs: Information systems Social networks
1. Introduction
@@
Social media has evolved into a vital arena for public discourse, where individuals and communities discuss political, social, and cultural issues. These platforms have fostered large-scale movements like the Arab Spring and Black Lives Matter, enabling activists and citizens to amplify calls for justice and change (Howard and Hussain, 2013; Freelon et al., 2016). However, social media has also been linked to polarization and radicalization, as seen in events surrounding the U.S. Capitol attack (Election Integrity Partnership, 2020; Luceri et al., 2021b), where online platforms played a role in mobilizing and coordinating participants  (Vishnuprasad et al., 2024). Similar cases include the Christchurch mosque shootings (Fakhrurroja et al., 2019) and far-right rallies across Europe (Klein and Muis, 2019; Pérez-Curiel, 2020), where social media content has exacerbated divisive sentiments and mobilized fringe communities (Davey and Ebner, 2019). The influence of online discourse has contributed to the establishment of state-sponsored information operations aimed at steering public opinion by introducing false or misleading narratives on popular platforms (Nisbet and Kamenchuk, 2019; Wilson et al., 2018; Molter and DiResta, 2020). Online platforms actively monitor and attempt to mitigate the integrity threats posed by information operations (TikTok, 2024; Meta, 2024; YouTube, 2024).
@@
Coordinated inauthentic behavior is a prominent tactic for distorting public discourse by amplifying specific viewpoints and giving the illusion of widespread support (Giglietto et al., 2020; Khaund et al., 2021; Pacheco et al., 2021; Murero, 2023; Cima et al., 2024). These activities typically involve synchronized posting and co-activity behavior patterns (Oates and Gray, 2019; Pacheco et al., 2021; Luceri et al., 2024b)— such as similar retweets, hashtag sequences, link sharing— designed to push particular narratives to the forefront of public attention. Studies have also documented how these coordinated behaviors contribute to detrimental social effects, such as the spread of propaganda (Hristakieva et al., 2022a), conspiracy theories (Tardelli et al., 2024a), and the promotion of disinformation (Weber and Neumann, 2021). Moreover, coordinated online actions have been linked to heightened toxicity in online conversations (Loru et al., 2024) and the dissemination of extremist ideologies (Kursuncu et al., 2021; Vishnuprasad et al., 2024).
@@
In recent years, research has increasingly examined social media’s role in political elections. Studies on Brazilian elections highlight how bots and coordinated networks shape public opinion, raising concerns on computational propaganda (Woolley and Howard, 2017; Pacheco, 2024). Similarly, research on European elections has shown how misinformation campaigns spread across social media can alter public perceptions, as seen in countries such as France, Germany, and Italy (Ferrara, 2017; Lovari, 2020; Baqir et al., 2024; Nogara et al., 2024). In the United States, investigations reveal how coordinated bot activity and foreign influence operations distorted online discussions and amplified divisive content across election cycles (Bessi and Ferrara, 2016; Howard et al., 2018; Luceri et al., 2019b; Tardelli et al., 2024a; Luceri et al., 2021a; Ferrara et al., 2020). However, most of this literature focuses on identifying such operations within individual platforms, predominantly Twitter, with limited investigation across the broader information space (Gatta et al., 2023).
@@
In this work, we expand on current research in coordination detection by broadening our scope to include the identification of coordinated inauthentic activity (CoIA) spanning multiple platforms. We adhere to the definition of “coordinated and inauthentic behavior” provided by Meta111about.fb.com/news/tag/coordinated-inauthentic-behavior/ and Twitter/$\mathbb{X}$222help.x.com/en/rules-and-policies/authenticity. We build on, adapt, and advance state-of-the-art coordination detection techniques to identify intra- and cross-platform CoIA that promotes external web domains and amplify specific narratives in the context of the 2024 U.S. Presidential Election. We leverage a large-scale dataset covering election-related online conversations spanning several platforms (Blas et al., 2024; memo6), including Twitter/$\mathbb{X}$, Facebook, and Telegram, collected during the run-up months to the Presidential Election. Combining insights from computational techniques for CoIA detection and language models for content analysis, we aim to answer the following Research Questions (RQs):
@@
Leveraging our multi-platform dataset and advancing CoIA detection techniques, we uncover multiple networks of coordinated accounts performing CoIA. We analyzed the textual content and link-sharing behaviors of these accounts, identifying the narratives they aim to amplify and the external domains they direct traffic to, assessing both their credibility and ties to foreign and domestic entities. Our findings reveal coordinated efforts to promote Russian-affiliated media across Telegram and $\mathbb{X}$, with highly partisan, low-credibility content systematically amplified by these networks. Conspiracy theories surrounding public health, the environment, and political topics, such as immigration and geopolitical tensions, were prominently featured. Notably, QAnon-related narratives were especially prevalent on Telegram, with coordinated accounts driving much of the discussion. Furthermore, we assessed the prevalence of AI-generated content produced by coordinated actors and the level of engagement their content attracted. We found that coordinated actors on Telegram relied on AI-generated content significantly more than organic users, while the opposite trend was observed on Facebook. This study sheds new light on cross-platform coordination efforts related to the upcoming U.S. Presidential Election, revealing the complex dynamics of influence operations. These findings underscore the urgent need for regulatory measures that go beyond individual platforms to effectively tackle the growing challenge of cross-platform coordinated inauthentic activity.
2. Related Work
@@
Influence Campaigns Interfering with U.S. Elections. A growing body of research has examined the influence of coordinated disinformation campaigns in U.S. elections, particularly efforts by foreign and domestic actors to manipulate social media during the 2016 and 2020 Presidential Elections (Sharma et al., 2022; Ferrara et al., 2020; Bessi and Ferrara, 2016; Vishnuprasad et al., 2024). For example, the Russian Internet Research Agency (IRA) was linked to a large-scale disinformation campaign during the 2016 election, deploying thousands of social bots and state-sponsored trolls to promote divisive narratives and foster discord (Badawy et al., 2019, 2018). The activity of these inauthentic actors has been extensively studied on Twitter (Zannettou et al., 2020; Luceri et al., 2020; Addawood et al., 2019; Luceri et al., 2019a), with fewer studies examining their presence on other platforms (Zannettou et al., 2019b, a).
@@
Similarly, the 2020 U.S. Presidential Election witnessed the propagation of false claims and conspiracy theories, including narratives about voter fraud and COVID-19 misinformation (Ferrara et al., 2020). Examples include the “Stop the Steal” movement, which spread across platforms like Twitter and Facebook, inciting allegations that the election results were fraudulent (Election Integrity Partnership, 2020; Vishnuprasad et al., 2024). Coordinated activity surrounding this narrative was widespread, often involving amplification techniques such as automated retweets or sharing similar content across accounts to create an illusion of widespread consensus (Luceri et al., 2021b).
@@
Detecting Coordinated Inauthentic Activity. Coordinated inauthentic activity has been identified as a predominant tactic in spreading disinformation and conspiratorial content (Luceri et al., 2024b; Pacheco et al., 2020). While not always malicious, coordination can also occur in legitimate organizing efforts, such as social movements. However, in disinformation campaigns, it is typically characterized by manipulative actions aimed at amplifying highly partisan content (Pacheco et al., 2021; Nizzoli et al., 2021).
@@
The detection of such deceptive, orchestrated efforts has evolved to address both automated and human-coordinated activities through advanced machine learning and network-based approaches. Traditionally, machine learning focused on identifying automated activities, such as bot-like behavior, and distinguishing them from human actions (Yang et al., 2019; Cresci et al., 2016). Recently, attention has shifted toward characterizing suspicious human-operated accounts, with research emphasizing content-, behavioral-, and sequence-based methods to detect coordinated actions by state-sponsored trolls (Hristakieva et al., 2022b; Alizadeh et al., 2020; Luceri et al., 2024a, 2020; Nwala et al., 2023; Ezzeddine et al., 2023).
@@
Network-based detection has gained prominence due to its ability to reveal coordinated behavior by constructing networks that highlight similarities in user actions, such as shared content, hashtags, or synchronized posting times (Pacheco et al., 2020, 2021; Nizzoli et al., 2021; Mannocci et al., 2024; Magelinski et al., 2022; Luceri et al., 2024b). Typical network-based metrics, such as centrality measures or edge weights, help identify clusters of users involved in coordinated Information Operations (IOs) within individual platforms, mainly Twitter (Luceri et al., 2024b; Vishnuprasad et al., 2024; Tardelli et al., 2024b).
@@
While previous work analyzed cross-platform harmful dynamics (Ng et al., 2022; Gatta et al., 2023; Zannettou et al., 2019a), our approach introduces a fine-grained unsupervised coordination detection framework by combining multiple network measures (edge weights and node centralities) to identify highly coordinated accounts engaged in CoIA across multiple platforms.
3. Data
@@
We collected data to analyze online discourse surrounding the 2024 U.S. Presidential Election across multiple social media platforms. Our data collection spans May and June 2024 and covers three major online platforms: Facebook, $\mathbb{X}$ (formerly Twitter), and Telegram. Each platform was queried for posts or messages containing specific election-related keywords (see Appendix A).
@@
Facebook data was collected through (the now defunct) Crowdtangle, which offered access to public posts of groups and pages. For $\mathbb{X}$, we gather publicly available information, including original tweets, retweets, replies, and quotes, retrieved via the platform’s web interface. Finally, Telegram data was gathered using the Telegram API, allowing the extraction of public chats’ details, metadata, messages, and message attachments. Details on the $\mathbb{X}$ and Telegram data collection infrastructure can be found in (Blas et al., 2024; Biswas et al., 2024).
To ensure consistency across all three platforms, we applied an identical filtering criterion, restricting the dataset to posts and messages published between May 1, 2024, and June 30, 2024.
Table 1 presents a summary of the datasets, including the number of accounts, [^3] We will use the term accounts to also refer to Facebook pages, Facebook groups, and Telegram channels. posts, and unique URLs for each platform.
@@
Platform Pages Accounts Channels Posts Tweets Messages URLs Domains
@@
Table 1. Dataset statistics across platforms.
4. Methodology
4.1. Exposing CoIA within and across Social Media Platforms
@@
Coordinated accounts employ various strategies to run campaigns; here, we analyze two key approaches: promoting specific URLs (or web domains) and generating highly similar textual content to amplify topics. The former creates the illusion of public consensus by artificially boosting links to external webpages, mock websites, and other social media networks (Minici et al., 2024; Gabriel et al., 2023). The latter manipulates platform feed algorithms by pushing specific keywords or hashtags, attempting to boost trending topics and inflate content popularity (Pacheco et al., 2021; Vishnuprasad et al., 2024). We refer to these orchestrated efforts as web domain promotion and content amplification, respectively.
Since URL-sharing follows a consistent mechanism across platforms, we analyze web domain promotion both within and across platforms, using unique URLs embedded in posts to track coordination. In contrast, textual similarity is analyzed only within individual platforms due to significant variability in content formats and constraints. Structural differences, such as character limits and messaging styles (e.g., Telegram chats vs. Facebook and $\mathbb{X}$ posts), make cross-platform textual analysis unreliable.
4.1.1. Detection of Web Domain Promotion
@@
To detect orchestrated campaigns promoting specific web domains, we analyze user similarities based on shared URLs, including both platform-specific and outbound links. To ensure data quality and reduce noise, we applied several preprocessing steps. First, we set a minimum activity threshold, requiring each user to share at least 10 unique URLs. This criterion, consistent with prior work (Luceri et al., 2024b; Pacheco et al., 2021), filters out low-engagement users who could alter the detection of coordinated accounts. Next, we expanded shortened or obfuscated URLs using a URL expansion library444github.com/dfreelon/unspooler, a common practice in the literature (Lasser et al., 2022).
@@
Co-URL similarity network. We constructed a bipartite user-URL network, where users connect to URLs they share. In the network’s adjacency matrix, rows correspond to users and columns to URLs. Following prior work (Luceri et al., 2024b; Pacheco et al., 2021), we applied Term Frequency-Inverse Document Frequency (TF-IDF) to represent the user-URL matrix. To prevent bias toward overly frequent URLs, we set a maximum “document” frequency at the 90th percentile and a minimum threshold of five occurrences per URL. This filters out both extremely rare and overly common URLs, ensuring a more meaningful set of URLs.
The bipartite user-URL network is transformed into a user-to-user similarity network by comparing pairwise co-shared URLs. Specifically, this similarity network is constructed by computing the pairwise cosine similarity between user TF-IDF vectors within the bipartite graph. This measure serves as the edge weight, quantifying the degree of similarity between users based on their shared URLs.
@@
Extraction of network properties for coordinated accounts detection. To identify online coordination, we build upon and integrate state-of-the-art strategies that detect CoIAs by filtering either low-weight edges or peripheral nodes from the similarity network. The first approach (Pacheco et al., 2021) emphasizes similarity strength, filtering out low-weight edges that likely represent spurious similarities. The second approach (Luceri et al., 2024b) focuses on the breadth of similarities, pruning nodes based on their centrality, i.e., the number of connections in a similarity network. This approach assumes that coordinated activity involves multiple accounts working together, which in a similarity network is expected to be reflected as a prominent collective similarity, with coordinated accounts linked to many well-connected nodes. As demonstrated by Luceri et al. (2024b), this makes eigenvector centrality an effective metric for detecting coordinated users.
Here, we extend this notion of collective similarity by considering the density of the similarity network. Network density quantifies how close a network is to full connectivity and, in this scenario, serves as a precise measure of collective similarity. Network density helps us combine node and edge filtering, balancing the breadth and strength of similarities for a more comprehensive coordination detection approach. This represents a conservative approach, which is especially important since these techniques have not yet been applied to platforms like Telegram and Facebook, where coordination dynamics may differ significantly.
Density-based unsupervised network dismantling. Our innovative approach leverages network density as a key variable to modulate the combination of node and edge filtering techniques. Specifically, edge weight and node eigenvector centrality are used jointly to identify groups of coordinated accounts by filtering out edges and nodes that fall below similarity and centrality thresholds, respectively. To achieve this in a fully unsupervised manner, we introduce a grid search method that systematically explores the two-dimensional parameter space of node centrality and edge similarity quantiles. The grid search evaluates network density under various threshold combinations, progressively filtering nodes and edges while measuring their impact on the density of connected components within the filtered similarity graph.
We hypothesize that as the filtering thresholds increase, a transitional phase in component density will emerge, signaling the presence of potentially coordinated accounts. To ensure robustness, we adopt a conservative approach, focusing on the minimum density among all connected components in the filtered similarity graph. By prioritizing the component with the lowest density, we ensure that all remaining components exhibit even higher densities, minimizing false positives—i.e., reducing the risk of misclassifying legitimate users as coordinated actors. Filtering thresholds are then selected based on the transitional phase of the lowest-density component, enabling the identification of highly suspicious CoIA. Details on the parameter selection are provided in the Appendix.
4.1.2. Detection of Content Amplification